OpenStudy (anonymous):

A number cube is rolled with these results: 64 ones, 67 twos, 73 threes, 59 fours, 72 fives, and 71 sixes. What is the experimental probability of rolling an even number? Write your answer as a percent, to the nearest tenth of a percent.

OpenStudy (anonymous):

64+67+73+59+72+71=406 Total Outcomes 67+59+71 = 197 even numbers probability of rolling an even number is \[\frac{197}{406}\] As a percentage multiply by 100 to give \[\frac{197}{406} \times 100=48.5%\]